OpenStudy (anonymous):

During a dry month, it rained three and two-thirds inches, which is nine and one-third inches less than the average. Consider the subtraction equation you would solve for the average monthly rainfall. What operation must be done on both sides to solve the equation?

OpenStudy (ranga):

Assume average rainfall is A inches. How would you represent "nine and one-third inches less than the average" ?

OpenStudy (anonymous):

Add nine and one-third.

OpenStudy (anonymous):

or subtract?

OpenStudy (ranga):

"less than the average" implies subtracting from the average. \[\Large A - 9\frac{ 1 }{ 3 } = 3\frac{ 2 }{ 3 }\]is the subtraction equation they are looking for. What operation must be done on both sides to solve for A?

OpenStudy (ranga):

You just need to add 9&1/3 to both sides to solve for A. A = 3&2/3 + 9&1/3 = (3 + 9) + (2/3 + 1/3) = 12 + 1 = 13

OpenStudy (ranga):

Or if you want it to change the mixed number into a fraction and then add it that would be okay too.\[A - 9\frac{ 1 }{ 3 } = 3\frac{ 2 }{ 3}\]\[A - \frac{ 28 }{ 3 } = \frac{ 11 }{ 3 }\]To both sides add \[\frac{ 28 }{ 3 }\]\[A = \frac{ 11 }{ 3 } + \frac{ 28 }{ 3 } = \frac{ 11 + 28 }{ 3 } = \frac{ 39 }{ 3 } = 13\]

OpenStudy (anonymous):

SO IT WOULD BE ADDITION I WAS RIGHT.